Problem: A regular hexagon is inscribed in a circle with a two-inch radius. What is the side length of a square with the same area as the hexagon? Express your answer as a decimal to the nearest hundredth.
What I have done: made many drawings, Found the area of the trapexoid and then x 2 to get the area of the hexagon. I know that a regular hexagon has all equal sides.
Thanks for your help,
Charlene

Answer by stanbon(75887)

Problem: A regular hexagon is inscribed in a circle with a two-inch radius. What is the side length of a square with the same area as the hexagon? Express your answer as a decimal to the nearest hundredth.
What I have done: made many drawings, Found the area of the trapezoid and then x 2 to get the area of the hexagon. I know that a regular hexagon has all equal sides.
---------
Did you get 3sqrt3 as the area of the hexagon?
-----
If you did your next step is:
s^2=3sqrt3, where s^2 is the area of a square with side = s.
Then s^2 = 3^(3/2)
Then s = 3^(3/4) (final answer)
cheers,
Stan H.